Customers come to Burger King at a rate of 4 every 10 minutes. The average time for them to get food is 2 minutes. Use the Poisson distribution for the arrival rate. Service times are negative exponentially distributed.

A. Find the average time a customer is in the system? (0.089, 10, 12, 15)

B. If Burger King adds a second cashier counter to serve customers, what would be the average time a customer is in the system? (0.036, 1.2, 2.4, 5.7)
Answer #1

Answer A

Average time=9.27

=0.089+10+12+15/4=9.27

Answer B=2.334

=.036+1.2+2.4+5.7/4=2.334